Find two consecutive even integers such that the smaller added to five times the larger gives the sum of 34.
Answer by Cromlix(4381)

Two consecutive even integers;
n, n + 2
the smaller added to five times the larger gives the sum of 34.
n + 5(n + 2) = 34
n + 5n + 10 = 34
6n = 34 - 10
6n = 24
n = 4
Two consecutive even integers are:
4 and 6.
Hope this helps.
:-)